Please show me how to solve the equation and check the answer. Thanks!
0.05(z+20)=0.1z-0.5

Answer by checkley71(8403)

.05(z+20)=.1z-.5
.05z+1=.1z-.5
.05z-.1z=-.5-1
-.05z=-1.5
z=-1.5/-.05
z=30
proof
.05(30+20)=.1*30-.5
.05*50=3-.5
2.5=2.5